TUGBOAT

Volume 1, Number 1 ^{/} October 1980

2 Editor’s Comments ^{/} Robert Welland

General Delivery 3 Message from the Chairman ^{/} Richard Palais 7 Publishing & TEX ^{/} Ellen Swanson 10 AMS-TEX — “A very friendly product” ^{/} Michael Spivak 12 Minutes of the first TUG meeting, February 1980 ^{/}

Robert Morris

Interface Software 16 The status of the Pascal implementation of TEX ^{/}

Ignacio Zabala and Luis Trabb-Pardo

17 The format of TEX’s DVI files ^{/} David Fuchs

Site Reports 19 University of Minnesota CDC Cyber site report ^{/}

Thea Hodge

Warnings & Limitations 20 Troubles with trace and Other oddities ^{/}

Barbara Beeton

Macros 20

Appendix A An indexing facility for TEX ^{/}

Terry Winograd and Bill Paxton

Questions & Answers 21

Letters 22 Gérard Emch and Arnold Pizer Miscellaneous Order form for The Joy of TEX Supplements TEX Errata

TUG mailing list

TUGBOAT

Volume 2, Number 1 ^{/} February 1981

2 Addresses of authors

General Delivery 3 Site Coordinators ^{/} Robert Welland 3 Chairman’s report ^{/} Richard Palais 5 Report on the January 1981 TUG Steering Committee meeting ^{/} Robert Morris 6 1980 TUG Treasurer’s report ^{/} Samuel B. Whidden 6 Informal TUG session ^{/} Robert Morris 7 A position on TEX maintenance ^{/} Robert Morris 9 TEX support ^{/} Samuel B. Whidden 10 Update on Pascal ^{/} Scott Kim

Interface Software 11 Erratum: The format of TEX’s DVI files ^{/} David Fuchs 11 TEX-Pascal and Pascal compilers (a status report) ^{/} Ignacio Zabala 12 TEX Font Metric files ^{/} David Fuchs 17 TEX support programs ^{/} Phil Sherrod and Alan Wright 20 Brief functional characterization of the procedures in the TEX/Pascal

compilation unit, SYSDEP ^{/} C. L. Lawson, I. Zabala and M. Díaz

32 Detailed specifications of procedures in the TEX/Pascal compilation unit, SYSDEP

/ C. L. Lawson, I. Zabala and M. Díaz

48 Output devices: A new column ^{/} Barry Doherty

Site Reports

· IBM Group 48 Status of TEX on the Stanford 370/3033 systems ^{/} Eagle Berns

· DEC 20 48 AMS site report ^{/} Barry Doherty and Barbara Beeton · VAX/UNIX 49 The status of VAX/TEX at Brown ^{/} Janet Incerpi · CDC Cyber 51 Report from the North Star, or, TEX at the University of Minnesota ^{/}

T. D. Hodge

· Univac 1100 51 TEX is available for Univac 1100 systems ^{/} Ralph Stromquist

· Nord 100 51 Report on the use of TEX at Computas A/S, Norway; Nord 100 computer: 16 bit

“big mini” ^{/} Helge Totland

Warnings & Limitations 53 Disappearing digits; Undisciplined uppercase ^{/} Barbara Beeton

Macros 53 How to prepare a file for publication in TUGboat ^{/} Barbara Beeton 55 TEX macro package ^{/} Max Díaz 56 Anatomy of a TEX macro package ^{/} Arthur M. Keller 87 NOFILL program with Pascal source ^{/} Lynne A. Price and Patrick Milligan 98 List macros ^{/} Lynne A. Price 111 Table of contents macros ^{/} Lynne A. Price 119 Utility macros ^{/} Patrick Milligan and Lynne A. Price 122 Hebrew letter (with source) ^{/} Lynne A. Price 125 Two slides ^{/} Lynne A. Price

Letters 127 Report from an early AMS-TEX user ^{/} David Eck

127 Graeme Hirst

Miscellaneous 128 Appendix A. Output samples from a paper by David Eck Supplements TEX Errata

TUG membership list

Volume 2, Number 2 ^{/} July 1981

2 Addresses of officers, authors and others 3 Official announcements

General Delivery 3 Editor’s remarks ^{/} Robert Welland 3 Chairman’s report ^{/} Michael Spivak 4 Report on the May 1981 TUG Steering Committee meeting ^{/} Robert Morris 5 TUG Treasurer’s report ^{/} Samuel B. Whidden 5 Proposal for institutional support of TUG ^{/} Robert Morris 6 Report: TEX Implementors’ Workshop, Stanford, 14–15 May 1981 ^{/} Barry Doherty 8 Workshop attendees 8 Preliminary announcement: TUG meeting, Cincinnati, January 1982 ^{/} Tom Pierce 9 Ask not what TUG can do for you, ask what you can do for TUG! ^{/} Patrick Milligan 10 A proposal for a machine independent tape interchange standard ^{/} Patrick Milligan

Software 12 The format of TEX’s DVI files, Version I ^{/} David Fuchs 16 Some feedback from PTEX installations ^{/} Ignacio Zabala 19 TEX and hyphenation ^{/} Frank M. Liang 21 TEX on small machines ^{/} Kent S. Harris and Robert M. McClure

Output Devices 25 Output device news flash (APS-5 and Linotron 202) ^{/} David Fuchs 25 Summary of computing equipment and output devices 26 Imagen (Canon LBP-10) ^{/} Luis Trabb-Pardo

Site Reports

· CDC Cyber 28 TEX under the North Star ^{/} Thea Hodge and Michael Frisch · DEC 10 & 20 28 DECSystem-10/20 Implementation Workshop announcement ^{/} Phil Sherrod

29 AMS site report ^{/} Barbara Beeton 29 TEX at the 1981 Spring DECUS U.S. Symposium ^{/} Patrick Milligan 29 TEX at NIH ^{/} Rachel Schwab

· Univac 1100 30 An implementation report for the Univac 1100 ^{/} Bill Kelly · VAX/VMS 33 Availability of Oregon Software implementation of TEX for the VAX/VMS ^{/} Monte Nichols

34 TEX for VAX/VMS ^{/} Barry Smith

Fonts 34 TUG Font Committee ^{/} Barry C. W. Doherty 35 The status of at OCLC ^{/} Thomas B. Hickey 39 Font development at the American Mathematical Society ^{/} William J. LeVeque 40 TPHON ^{/} Ronald Whitney 40 Proofmode and magnification ^{/} Barry Doherty and Ronald Whitney

Warnings & Limitations 42 Uppercase update; Fickle fonts ^{/} Barbara Beeton

Macros 43 Greetings from the Editor ^{/} Lynne Price 43 Macros on microfiche 43 Erratum: NOFILL program ^{/} Patrick Milligan 44 TEX macros for automatic font code allocation ^{/} Patrick Milligan 46 A macro menagerie ^{/} Brendan McKay 50 Macro madness ^{/} Michael Spivak

Appendix A Fácil TEX ^{/} Max Díaz

Problems 54 Greetings from the Editor ^{/} Lynne Price 54 Problems from the TEXarcana course 57 Balancing columns of text and translation ^{/} Johnny Stovall 57 Input-dependent macro redefinition ^{/} Michael Spivak

Letters 57 Lynne Price

Dreamboat 58 Lynne Price

Miscellaneous Instructions for submitting TUGboat articles on magnetic tape Membership application and order form

Supplements  TEX and  Errata; TUG membership list

Volume 2, Number 3 ^{/} November 1981

2 Addresses of officers, authors and others 3 Official announcements

General Delivery 3 Message from the Chairman ^{/} Michael Spivak 4 TUG Treasurer’s report ^{/} Samuel B. Whidden 5 The current state of things ^{/} Donald Knuth 7 TEX Users Group Winter 1982 meeting, January 11–12, 1982, Cincinnati, Ohio ^{/}

Tom Pierce

Software 7 Pascal-coded TEX errata ^{/} Arthur Samuel 8 The format of PXL files ^{/} David Fuchs

Output Devices 13 Sample of output from an IDS-640 ^{/} John Sauter 14 Output device index ^{/} Rilla Thedford 14 A Varian output driver in VAX/VMS Fortran ^{/} Jim Mooney 15 Diabolic TEX ^{/} Timothy Murphy

Site Reports 21 News from the home front (Stanford) ^{/} David Fuchs

· CDC Cyber 22 TEX under the North Star ^{/} Michael Frisch · DEC 20 23 A TUGboat tour: Excerpts from the TEXnician’s log ^{/} Barbara Beeton · HP 3000 25 TEX for the HP3000 ^{/} Lance Carnes

· IBM Group 26 TEX for the IBM 370 ^{/} Susan Plass 27 TEX in Israel ^{/} Jeffrey S. Rosenschein 28 TEX at the University of Michigan, summary of progress ^{/} David Rodgers

· VAX/UNIX 29 VAX on UNIX ^{/} Robert Morris 29 TEX at CalTech ^{/} Calvin Jackson 32 The status of VAX/TEX at Brown ^{/} Janet Incerpi

· VAX/VMS 32 VAX/VMS site report ^{/} Monte C. Nichols 32 Enhancements of VAX/VMS TEX at Calma ^{/} John Blair 34 “Poor man’s” TEX ^{/} John Sauter

Fonts 35 Font update ^{/} Ronald Whitney

Warnings & Limitations 35 Don’t just \let TEX hang, \raise or \lower it ^{/} Barbara Beeton

Macros 36 Editor’s introduction ^{/} Lynne Price 36 TUGboat macro index 37 Bubbles: A TEXtension in search of a TEXpert ^{/} Timothy Murphy 39 Charting your grammar with TEX ^{/} Michael F. Plass 57 Chemical notation using TEX ^{/} Monte Nichols, Barbara Beeton

Problems 58 Lynne Price Balancing columns of text and translation (Johnny Stovall) Input-dependent macro redefinition (Mike Spivak) A macro that prints its name (Anthony Siegman) Testing the equivalence of strings

61 Problems from the TEXarcana class: Answers, and another problem

Letters et alia 66 J. M. Cole 66 Formatting a book with TEX: Experiences and observations ^{/} Michael Sannella

Miscellaneous Instructions for submitting TUGboat articles on magnetic tape Membership application and order form

Supplements  TEX and  Errata

TUG membership list

Volume 3, Number 1 ^{/} March 1982

2 Addresses of officers, authors and others 3 Official announcements

General Delivery 3 Statement of editorial policy ^{/} Barbara Beeton

4 Report on the January 1981 TUG Steering Committee meeting ^{/} Robert Morris

TUG Winter meeting, January 11–12, 1982, Cincinnati, Ohio 5 Program 6 Attendees 7 Open letter to TUG ^{/} Robert McGaffey and Keith Penny 10 Report on the ANSI X3J6 meeting ^{/} Lynne A. Price

Software 10 Fixed-point glue setting — an example of WEB ^{/} Donald Knuth

Site Reports

· CDC Cyber 28 CDC TEX ^{/} Michael J. Frisch

· HP 3000 28 HP3000 site report ^{/} Lance Carnes · IBM Group 29 TEX at the University of Milan ^{/} G. Canzii, D. Lucarella and A. Pilenga

30 TEX news from Pisa ^{/} L. Aiello and S. Pavan

· VAX/VMS 32 VAX/VMS site report ^{/} Monte C. Nichols

“small” TEX 32 Editor’s introduction ^{/} Lance Carnes

Warnings & Limitations 33 Another hangup ^{/} Barbara Beeton

Macros 33 Editor’s introduction ^{/} Lynne Price 33 TUGboat macro index 35 Display of a font in table form ^{/} Roger L. Beeman 39 Seating charts ^{/} Roger L. Beeman

Problems 43 Problems from the TEXarcana class: Answer to the challenge problem

Late-Breaking News TUG financial reports ^{/} Samuel B. Whidden 45 TUG Treasurer’s report 46 1982 TUG Budget

Miscellaneous Instructions for submitting TUGboat articles on magnetic tape Membership application and order form

Supplements  TEX and  Errata

TUG membership list

Volume 3, Number 2 ^{/} October 1982

2 Addresses of officers, authors and others 3 Official announcements

General Delivery 4 The TEX logo: An important note

4 Report on business meetings, TUG Summer meeting,

Stanford University, July 25–27, 1982 ^{/} Susan Plass

TUG Summer meeting and TEX82 short course,

Stanford University, July 25–30, 1982

5 Program 5 Attendees 7 An informal interchange format for TEX files ^{/} Pierre A. MacKay 9 Introduction to TEX and TUG for new users ^{/} Ron Whitney

Software 13 TEX82 memory structure 14 The format of TEX’s DVI files ^{/} David Fuchs

Site Reports 20 News from Stanford ^{/} David Fuchs

· IBM Group 21 Fixes to known bugs in TEX370 ^{/} Susan Plass

22 TEX installation at the University of Michigan ^{/} Paul Grosso

· VAX/VMS 23 VAX/VMS site report ^{/} Monte C. Nichols and David Kellerman

“small” TEX 24 Editor’s introduction ^{/} Lance Carnes

Fonts 25 A Fortran version of ^{/} Sao Khai Mong

Warnings & Limitations 25 Charting the generation gulf ^{/} Barbara Beeton

Macros 26 Font codes in popular use ^{/} Calvin Jackson 27 Editor’s introduction ^{/} Lynne Price 27 TUGboat macro index 28 Multi-column output format ^{/} Barbara Beeton 34 Some TEX programming hacks ^{/} Leslie Lamport 34 Unblocking an AMS-TEX tape ^{/} Barbara Beeton

Problems 38 Paragraphs in tables ^{/} Mark Blanford 38 Hanging punctuation

Late-Breaking News TUG financial reports ^{/} Samuel B. Whidden 39 TUG Treasurer’s report 40 1983 TUG Budget

Miscellaneous Membership application and order form

Supplements  TEX and  Errata

TUG membership list

Volume 4, Number 1 ^{/} April 1983

2 Addresses of officers, authors and others 3 Official announcements

General Delivery 4 Library subscriptions — What are they?

4 Users’ Course in AMS-TEX ^{/} Michael Spivak

TUG financial reports ^{/} Samuel B. Whidden 4 TUG Treasurer’s report 5 1983 TUG budget

Software 6 TEXhax summary ^{/} David Fuchs

Output Devices 10 Chart: Output devices and computers 10 Index to sample output from various devices 11 Low-cost downloadable font devices ^{/} Nelson H. F. Beebe 12 TEX on the OSP130 ^{/} Patrick Ion, Bill Hall, Rilla J. Thedford 13 Sample output from Florida Data OSP 130

Site Reports 14 News from all over ^{/} David Fuchs

· HP 1000 16 TEX on the HP-1000 ^{/} Irene J. Bunner and John D. Johnson · VAX/UNIX 17 Unix TEX site report ^{/} Richard Furuta and Pierre MacKay

18 Porting TEX to VAX/UNIX ^{/} Pavel Curtis and Howard Trickey 21 TEX at the University of Washington: Tops-20, Unix, Versatec,

and the Monolithic ^{/} Pierre MacKay and Richard Furuta

· VAX/VMS 22 VAX/VMS site report ^{/} Monte C. Nichols

23 TEX at Calma R & D ^{/} David Krapp

· DG MV8000 23 TEX at Texas A & M University ^{/} Norman Naugle and Bart Childs

“small” TEX 24 Editor’s introduction ^{/} Lance Carnes

Fonts 25 Pictures are just big letters ^{/} Scott Guthery 26 Computer calligraphy ^{/} Georgia K. M. Tobin 33 Announcement: Fifth ATypI Working Seminar,

The computer and the hand in type design:

The aesthetics and technology of digital letterforms

Macros 33 TUGboat macro index 35 How to build a \strut ^{/} Barbara Beeton 36 Determining hashtable size and other quantities ^{/} Barbara Beeton 37 Some layout macros ^{/} August Mohr 38 Testing the widths of a font ^{/} Robert M. McClure

Problems 39 Hanging punctuation

Letters et alia 40 How to obtain TEX82 on tape ^{/} Maria Code 40 Announcement: Manipulation de Documents — Journées Francophones

Dreamboat 41 TEX as a programming language? ^{/} A. E. Siegman

Advertisements 43 Textset, Inc. — A service for TEX users 44 Quality Micro Systems — Lasergrafix 1200

Miscellaneous 45 TEX82 order form 47 Membership application and order form

Supplements  TEX and  : Errata and changes

TUG membership list

Volume 4, Number 2 ^{/} September 1983

50 Addresses of officers, authors and others 51 Official announcements

General Delivery 52 Message from the President ^{/} Pierre MacKay

Minutes of the July 1983 TUG Steering Committee meeting 53 Susan Plass. July 13–14 54 Chuck Dupree. July 15 55 Summary of the technical program, July 1983 TUG meeting 57 Participants, TUG meeting and AMS-TEX82 short course 59 Final report of the Bylaws Committee ^{/} Susan Plass 60 TEX Users Group — TUG Operating Procedures 62 TUG Treasurer’s report ^{/} Samuel B. Whidden 63 Report on ANSI X3J6 ^{/} Lynne Price

Software 64 A note on hyphenation ^{/} Donald Knuth 65 TEX vs. INITEX ^{/} David Fuchs 66 TEXhax summary ^{/} David Fuchs

Output Devices 71 Chart: Output devices and computers

Site Reports 72 News from the TEX Project ^{/} David Fuchs · Honeywell CP-6 73 TEX82 on CP-6 ^{/} Rick Mallett

· IBM Group 74 IBM site report ^{/} Susan Plass

· UNIX 74 Unix TEX site report ^{/} Richard Furuta · VAX/VMS 75 VAX/VMS site report ^{/} Monte C. Nichols

76 Revised Varian output driver in VAX/VMS Fortran ^{/} Jim Mooney

Fonts 76 TEX for Arabic script ^{/} Pierre MacKay

“small” TEX 77 Table of “small” TEX implementations ^{/} Lance Carnes

Macros 78 Problems from the TUG meeting:

Framed slides; Multiple marks ^{/} Lynne Price

79 TUGboat macro index

Problems 80 First line of paragraph ^{/} Jim Sterken

Letters et alia Observations on TEX from a divergent viewpoint 81 Sam Whidden. Introduction 82 J. R. Roesser. Memo to STI staff:

TEX Users Group meeting, July 1983 Comments and responses

90 Don Knuth 90 David Fuchs 93 Michael Spivak 95 Richard Palais 97 Barbara Beeton 99 J. R. Roesser. Response to the responses 102 Barbara Beeton. Wrapup

Late-Breaking News 103 Summary of AMS-TEX ^{/} Michael Spivak

Advertisements 127 Textset, Inc. 127 TEX lectures on tape 128 Imagen Corporation — Intelligent page printer systems

Miscellaneous 129 TEX82 Order Form 131 Membership Application and Order Form

Supplements  TEX and  : Errata and changes

TUG membership list

Volume 5, Number 1 ^{/} May 1984

2 Addresses of officers, authors and others 3 Official announcements

General Delivery 4 Message from the President ^{/} Pierre MacKay 4 TEX incunabula ^{/} Donald E. Knuth. 12 TUG 1983 financial report with comparisons ^{/} Samuel B. Whidden. 13 Preliminary program — August 1984 Stanford meeting

Software 14 Hyphenation of Italian words ^{/} G. Canzii, F. Genolini and D. Lucarella 15 Hyphenation exception log

Output Devices 16 Chart: Output devices and computers 17 Index to sample output from various devices 17 How to tame your phototypesetter by TEX ^{/} Goffredo Haus 18 DVI-code to the Harris 7500 ^{/} Han Noot

Site Reports 22 News from the TEX Project ^{/} David Fuchs · CDC Cyber 23 CDC TEX at RECAU ^{/} Benedict Løfstedt · IBM Group 24 TEX at Washington State University ^{/} Dean Guenther,

Alan Hagen-Wittbecker and Janene Winter

· Data General 26 Data General site report ^{/} Bart Childs · Prime 26 Prime site report ^{/} Bart Childs · UNIX 27 Unix TEX site report ^{/} Richard Furuta

28 TEX 1.0 on Sun workstations ^{/} D. L. Rodgers, J. J. Sterken and P. Grosso

· VAX/VMS 29 VAX/VMS site report ^{/} Monte C. Nichols 30 Notice to VAX/VMS users ^{/} David Kellerman and Barry Smith

Fonts 31 Metafont Generic Font file format 34 Gray fonts for Metafont proofs 36 The OCLC Roman family of fonts ^{/} Georgia K. M. Tobin

“small” TEX 47 Table of “small” TEX implementations ^{/} Lance Carnes

Warnings & Limitations 48 \relax and watch the numbers ^{/} Barbara Beeton

Macros 49 The HP TEX macros ^{/} Susan Daniels

Letters et alia 67 Comments on quality in publishing ^{/} Don Knuth

Advertisements 67 Classified 68 Textset, Inc.

Miscellaneous 69 TEX82 Order Form 71 Membership application and order form

Supplements The TEXbook: Errata and changes TUG membership list

Volume 5, Number 2 ^{/} November 1984

74 Addresses of officers, authors and others 75 Official announcements

General Delivery 76 Message from the President ^{/} Pierre MacKay 77 Report of the Publications Committee ^{/} Robert Welland 77 Report of the Special Projects Committee ^{/} Arthur Keller 78 Submitting items to TUGboat ^{/} Barbara Beeton 79 First principles of typographic design for document production ^{/}

Richard Southall

Software 91 Editor’s introduction ^{/} Helmut Jürgensen

TEXtensions 91 How to run TEX in a French environment:

Hyphenation, fonts, typography ^{/} Jacques Désarménien

103 German hyphenation and Umlauts in TEX ^{/} Bernd Schulze

Fonts 105 A course on programming ^{/} Donald E. Knuth 119 A Chinese meta-font ^{/} John D. Hobby and Gu Guoan

Output Devices 137 Chart: Output devices and computers 138 Index to sample output from various devices 138 Dvi QMS: an example of a driver ^{/} Tom Rokicki 139 \special ^{/} Tom Rokicki

Site Reports 141 News from the TEX Project ^{/} David Fuchs

141 TEX users’ activity in Germany · Prime 141 Prime site report ^{/} John Crawford · UNIX 142 Unix TEX site report ^{/} Richard Furuta

· VAX/VMS 142 VAX/VMS site report ^{/} Barry Smith 143 Feedback from TEX users at Lockheed ^{/} Carlos A. Felippa

“small” TEX 145 Table of “small” TEX implementations ^{/} Lance Carnes

Letters 146 Mathematical communication with a deaf and blind student using TEX ^{/} Malcolm Clark

News & Announcements 146 Calendar

147 Call for papers: TEX for scientific documentation, Varenna, Italy, May 16–17, 1985

TUG Business 148 TEX Users Group 1984 financial report with comparisons ^{/} Samuel B. Whidden 149 Report of the August 1984 Steering Committee meeting ^{/} Chuck Dupree 153 Profile of TEX installations available to TUG members

Advertisements 156 Talaris talks TEX 158 Quality Micro Systems, Inc. 159 Textset, Inc. 160 The Metafoundry 160 Instructors wanted for TUG courses

Miscellaneous 161 Membership application and order form 165 Order form for American Mathematical Society publications 167 TEX82 order form

Supplements The TEXbook: Errata and changes TEX Users Group Membership list